• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Humensky

    Log entry time 22:39:52 on July 18,1999

    Entry number 23078

    keyword=status update

    When swing shift began, the accelerator was still recovering from the CHL crash. Around 6pm beam returned and it was quickly evident that the tune was still improper... see figure 1. The large amplitude of oscillation on target is unacceptable, as is the identical response of the 3a and 3b monitors to all the coils. The crew performed some tests, several requested by the optics on call Chao, but thunderstorms precluded checking the results of the last changes they made. Shortly after 8pm the thunderstorm caused several water leaks on magnets and tripped the safety interlocks, sending the north linac to restricted access. At that time, MCC predicted at least several hours of downtime.